Please help me solve this question by using the Quadratic formula to
-x^2 - 3x = -5 for x.
Thank you

Answer by TimothyLamb(4379)

-x^2 - 3x = -5
-x^2 - 3x + 5 = 0
---
the above quadratic equation is in standard form, with a=-1, b=-3, and c=5
---
to solve the quadratic equation, plug this:
-1 -3 5
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
the roots of the quadratic are:
x = -4.1925824
x = 1.1925824
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php